An engine approaches the hill with a constant speed. When it is at a distance of 0.9 km, it blows a whistle whose echo is heard after 5 seconds. What is the speed of engine. Speed of sound is 330 m/s

Grade:12th pass

2 Answers

Shashwat Avasthi
8 years ago
Let speed of train be v. The echo is heard after the sound travels to the hill and then back. Original distance=900m 
Distance travelled by train in the 5s before echo=5v
So distance by sound in 5s= 1800-5v
330=360-v; v=30m/s
Vignesh
7 years ago
Let speed of train be v. The echo is heard after the sound travels to the hill and then back. Original distance=900m Distance travelled by train in the 5s before echo=5vIf we apply Doppler effect, 1800 - 5v=(330+v)5 360 -v=330+v v=15m/s.. Why is this wrong??